OpenStudy (anonymous):

. If the two angles below are supplementary, solve for x. (Points : 2) 180 156 12 13

OpenStudy (anonymous):

@Brittni0605 @cherio12 I GOT 12 AM I RIGHT?

OpenStudy (anonymous):

13

OpenStudy (anonymous):

180=7x+5x+24 180=24+12x

OpenStudy (anonymous):

180=36X @cherio12

OpenStudy (anonymous):

UGHH

OpenStudy (anonymous):

no, you can't combine 24 and 12. You can only combine like terms

OpenStudy (anonymous):

180-24=12x 156=12x x=13

OpenStudy (anonymous):

THAN WHATS MY ANSWER?

OpenStudy (anonymous):

13